OpenStudy (anonymous):

decompose x^4 - x^3 - 2x^2 + 4x - 3 / x(x-1)^2 into partial fractions.

OpenStudy (lgbasallote):

hint: your denominator is x(x-1)^2 so your PF. would look something like \[\frac Ax + \frac B{x-1} + \frac C{(x-1)^2}\] does that help?

OpenStudy (anonymous):

hmmm... you have to divide first, because the degree of the denominator is less than the degree of the numerator

OpenStudy (lgbasallote):

lol..didn't see that

OpenStudy (anonymous):

so I must first do long division?

OpenStudy (anonymous):

yes, this problem pretty much sucks, because there is no avoiding the long division i would cheat http://www.wolframalpha.com/input/?i=%28+x^4+-+x^3+-+2x^2+%2B+4x+-+3+%29%2F+%28x%28x-1%29^2%29

OpenStudy (anonymous):

How do i know that in this case the degree of the the Denominator is more than the degree of the numerator?